# Modular Compilation for Quantum Chiplet Architectures

### v2

@@ Abstract.
As quantum computing technology continues to mature, industry is adopting modular quantum architectures to keep quantum scaling on the projected path and meet performance targets. However, the complexity of chiplet-based quantum devices, coupled with their growing size, presents an imminent scalability challenge for quantum compilation. Contemporary compilation methods are not well-suited to chiplet architectures — in particular, existing qubit allocation methods are often unable to contend with inter-chiplet links, which don’t necessary support a universal basis gate set. Furthermore, existing methods of logical-to-physical qubit placement, swap insertion (routing), unitary synthesis, and/or optimization are typically not designed for qubit links of significantly varying levels of duration or fidelity. In this work, we propose SEQC, a complete and parallelized compilation pipeline optimized for chiplet-based quantum computers, including several novel methods for qubit placement, qubit routing, and circuit optimization. SEQC attains up to a 36% increase in circuit fidelity, accompanied by execution time improvements of up to 1.92×1.92$\times$1.92 ×. Additionally, owing to its ability to parallelize compilation, SEQC achieves consistent solve time improvements of 24×2-4$\times$2 - 4 × over a chiplet-aware Qiskit baseline.
@@ 1. Introduction
We are already observing signs of this shift. Recent developments in quantum chip linking, including flip-chip architectures (gold2021entanglement) and low-loss coaxial cables (niu2023low), suggest that modular designs are the most viable for scaling quantum computers (smith_2022_chiplets). Many contemporary or upcoming leading quantum systems adopt chiplet-based modular quantum processors, for example using high-bandwidth quantum links between nearest-neighbor quantum chiplets (e.g., carrier-chip couplers in Rigetti Aspen-M (Rigetti_AspenM; gold2021entanglement), or m-couplers in IBM Crossbill (gambetta_2024_IBMcouplers; Mandelbaum_2024_IBMCouplers)), or lower-fidelity, lower-bandwidth, but longer-distance flexible coupling of discrete chips (e.g., l-couplers in IBM Flamingo (gambetta_2024_IBMcouplers; Mandelbaum_2024_IBMCouplers)), or a combination of the above (e.g., c-, m-, and l-couplers in IBM Starling (Mandelbaum_2024_IBMCouplers)), or multi-chip connectivity through tunable couplers and routing chips (Field_2024_RigettiModular). Most major companies have set their sights at modular designs to meet quantum scaling targets in practical ways, and modular quantum processors dominate their latest roadmaps (IBM_2023_Roadmap; IonQ_2024_Roadmap; Rigetti_2024_Roadmap).
@@
Unfortunately, current compilation infrastructure is not capable of reasoning well about this quantum interconnect heterogeneity. The mapping of a quantum program’s logical quantum gates into the native gates supported by the underlying quantum processor, and the mapping of logical qubits to the physical qubits of the quantum hardware, largely determine the number and type of native quantum gates required for the computation, the circuit depth, and its execution time—long programs may not complete successfully as qubits decohere. Different mappings can result in vastly different compiled quantum circuits, with diverse characteristics along all these axes. These details of efficiency can make or break an algorithm in today’s noisy intermediate-scale quantum systems (NISQ) era, so quantum compilers aggressively optimize for all of them. To make matters more challenging, each qubit, coupler, and gate have diverse error profiles that are highly variable both spatially and temporally (Tannu_2019_ErrorVariability; Murali_2019_NoiseMapping; dasgupta2021stability). Unlike classical compilation that is done only once for a given architecture, quantum programs must be recompiled every time before execution, as the ideal physical qubits to execute on change between runs. So, quantum compilers today are faced with the daunting task of optimizing quantum programs across multiple dimensions with often conflicting demands, in a continuously changing environment. Coupled with the hardness of synthesizing unitaries (which is exponential in the number of qubits) and the need for quick and frequent compilation, modern quantum compilers have no option other than to rely on heuristics to perform the task. These heuristics result in a compilation complexity that we experimentally observed to be $O(n^{2})$ for $n$ total qubits in a quantum processor, as shown in Figure 8.
@@
Compilation in classical systems is typically performed independently and in parallel for each source file, producing one object file per source. The individual object files are then linked together to construct the executable. We propose a compilation framework for modular quantum processors that works in a similar fashion: it stratifies, i.e., splits, the source quantum circuit into subcircuits small enough to fit in each chiplet, and maps subcircuits to chiplets, and then in parallel elaborates each subcircuit and compiles it for its target chiplet. This Stratify-Elaborate Quantum Compiler (SEQC) stratifies a source program only once for a given chiplet architecture, and performs only the elaboration step recurrently before each execution. In essence, SEQC replaces the recurrent $O(n^{2})$ compilation step for an $n$-qubit quantum processor, with several parallel $O(k^{2})$ elaboration steps for $k$-qubit chiplets. As the qubit capacity $n$ of quantum processors grows exponentially, today’s $O(n^{2})$ compilation latency rises even faster. We expect, however, that the number of qubits $k$ per chiplet will remain relatively stable or grow much slower, as it seems to be the case for the foreseeable future (IBM_2023_Roadmap), and thus the SEQC recurrent compilation latency is expected to remain relatively stable. The stratification step is $O(n^{2})$, so the end-to-end complexity remains the same, but stratification is performed only once; the recurrent compilation in SEQC is only $O(k^{2})$, and barely growing with new processor designs.
Additionally, as SEQC is cognizant of hardware modularity, it can stratify the source quantum circuit into subcircuits and map them to chiplets to minimize inter-chiplet communication. As we show in this paper, SEQC produces circuits with shorter execution times and significantly fewer inter-chiplet gates compared to today’s stock compilers, leading to much higher fidelity execution. More importantly, as the number of qubits in a processor grows, SEQC achieves even higher performance in these figures of merit.
@@
  • We design and implement in SEQC several novel methods for qubit placement, qubit routing, and circuit optimization.
  • We evaluate SEQC and show it compiles circuits with up to 36% higher circuit fidelity and up to 1.92×1.92$\times$1.92 × lower execution time, while consistently achieving 24×2-4$\times$2 - 4 × faster compilation time compared to a chiplet-aware Qiskit baseline.
@@ 2. Background and Motivation
(a) A chiplet quantum architecture
(b) Input circuit from SABRE
@@ 2.1. Qubit Allocation
Qubit allocation addresses the topology constraints of quantum devices, where physical qubits on a device have limited connectivity to other qubits on the device. It is critical that after this stage, which often requires the modification or addition of circuit gates, the resulting optimized and technology-dependent circuit is functionally equivalent to the original, technology-independent algorithm (smith2019quantum). The qubit allocation problem has been previously shown to be NP-complete (qubit_allocation); thus, algorithms often adopt various strategies to make the problem more tractable. One common method is to further divide qubit allocation into two substages: layout and routing.
Logical-to-physical qubit layout, also known as qubit placement, involves deriving an injective map from the virtual qubits of the quantum circuit to the physical qubits of the quantum device (qubit_routing; sabre_swap). In other words, the qubit layout stage decides the initial position of logical qubits on the device. In contrast, qubit routing operates from an initial qubit layout and inserts SWAP operations to align a quantum circuit’s gates to a given qubit topology. Consider, for example, the state-of-the-art in qubit allocation: the SABRE heuristic search algorithm (sabre_swap). In SABRE, the distinction of the layout and routing stages motivates an iterative method for solving qubit allocation. Namely, an initial layout is used to perform qubit routing, which is used to generate a new initial layout.
@@
Because qubit routing remains NP-hard (qubit_routing), another method of simplifying the problem is to assume qubit links are roughly equivalent, particularly with respect to average fidelity and two-qubit gate duration. This assumption can manifest implicitly when a given device topology is represented as an unweighted, undirected graph. For monolithic architectures, one could debate the marginal tradeoff between performance and solution quality; however, for modular architectures, the cross-chiplet connections are far worse compared to intra-chiplet connections (smith_2022_chiplets). Thus, we could expect the solutions generated from the simplified model to be worse than a more physically accurate model, for example, in circumstances where the shortest path in terms of the quantity or depth of inserted SWAPs diverges from the shortest path in terms of fidelity, gate duration, or some combination thereof.
2.2. Basis Translation
@@
Basis translation converts the gates of an input quantum circuit into gates that are physically supported on hardware. Monolithic quantum architectures typically support a so-called universal basis gate set, where any unitary quantum operation can be constructed using solely the gates in the set (mike_and_ike). Moreover, it has been shown that basis translation can be performed efficiently given the target gate set is universal, and the quantum circuit contains a fixed, finite number of qubits (Solovay–Kitaev). Unfortunately, modular quantum architectures depend upon specialized links to connect otherwise-independent chiplets. In the general case, the cross-chiplet links can only be used for data movement (such as SWAPs) and not computation, meaning these links may not support universal operations (niu2023low). Thus, compilers that depend on the expectation of universality among qubit links (qiskit_basis_translation_errors) and do not expect these heterogenous, non-universal basis gate sets can generate incompatible outputs when targeting modular architectures.
2.3. Optimization
@@
Quantum circuit optimization seeks to prune extraneous operations by finding opportunities to combine, eliminate, and/or parallelize quantum gates (optimization1). Through commutation and cancellation identities, it is possible to search for optimization opportunities by iteratively transforming the quantum circuit without affecting the underlying operation (optimization1). However, in general, optimization is highly time- and resource-intensive, with the problem scaling exponentially with respect to number of qubits and/or circuit depth (qfactor). As a result, partitioning, the division of a quantum circuit into blocks of quantum gates, is a common tactic among optimization techniques (qfactor; qgo; bqskit). One major consideration in this process is demarcation, i.e., where and when to draw the line between blocks. Namely, one forfeits all optimization opportunities along boundaries of demarcation, so they must be carefully chosen to minimize negative impact on the output solution. From this perspective, modular architectures are naturally synergistic with partitioning, since cross-chiplet connections present a physically motivated boundary between operations.
3. A Strawman Chiplet Compiler
@@
As discussed previously in Section 2, our model of modular quantum architectures includes the more general case where inter-chiplet links are can be non-universal. This is representative of hardware manufacturing constraints (niu2023low), where only a restricted set of operations, e.g., state transfers or SWAPs, may be supported over long-distance couplers (inter-chiplet links). However, many compiler designs (including designs for modular architectures, see Section 7) experience errors or crashes when presented with non-universal, heterogenous, basis gate sets. For example, the Qiskit compiler (qiskit2024) is explicit in its compatibility for only universal basis gate sets (qiskit_basis_translation_errors). While it is possible to create a Qiskit Backend that supports heterogeneous basis gate sets, the compiler itself does not take them into account. Specifically, the Qiskit routing stage (zou2024lightsabrelightweightenhancedsabre) (an implementation of SABRE) will freely place gates on incompatible physical links, and there are no methods in the compiler to prevent or fix consequent errors. For the purposes of this work, however, it was critical to establish a baseline representative of the current state of quantum compilation, against which we can compare our more-optimized design (see Section 4). Thus, we created a “Strawman” baseline compiler, built on top of the Qiskit compiler (qiskit2024), with the minimal modifications necessary to generate valid outputs for chiplet architectures.
@@
To accomplish this, we leverage a “peephole” compilation pass, which is commonly used in classical compilation to perform optimization (dragon_book). Peephole passes operate within a nominal peephole window and substitute its contents with a logically-equivalent replacement. Notable examples in the context of quantum compilation include basis translation and elimination of self-adjoint gates. This design strategy facilitates modular compilation, not only in the sense that it can be easily parallelized, but also because peephole passes can be added between other compilation passes without breaking compatibility. These qualities make a peephole pass well-suited to the Strawman baseline, since we do not want to make extensive modifications to any of Qiskit’s built-in compilation stages or passes, nor introduce a significant performance overhead. The peephole pass used in the Strawman compiler is inserted between the routing and translation compilation stages in Qiskit. As shown in Figure 1, it operates on the circuit generated by SABRE (zou2024lightsabrelightweightenhancedsabre), where it detects when non-SWAP gates are placed on non-universal inter-chiplet links and employs a cyclic qubit rotation, i.e., a perfect shuffle permutation, to move the aforementioned gates to nearby intra-chiplet links ( see Figure 1(c)). Overall, for every inter-chiplet gate in the original circuit generated by SABRE, the peephole pass inserts 4 additional SWAP gates—2 inter-chiplet and 2 intra-chiplet. Ultimately, this degree of overhead is inevitable when the stock, unmodified compiler is ignorant of chiplet-specific device constraints.
@@
To emphasize the rigor of the Strawman, it is worth discussing the designs that were ultimately discarded. For example, consider a hypothetical alternative design that skips the final peephole step shown in Figure 1(d). This would appear to save $2$ SWAPs, thus halving the cost per inter-chiplet gate. However, the hypothetical pass would also no longer be a true peephole, since it introduces a net qubit permutation that propagates to downstream gates in the circuit. From the perspective of modular compilation, this results in serial dependencies that prevent parallel execution of the pass. From a routing perspective, the resultant qubit permutations disrupt the solution generated by SABRE, leading to destructive routing errors that can only be resolved by an entirely new routing pass. Overall, such a hypothetical pass results in substantially worse output circuits compared to the peephole method, which we confirmed through testing. Other potential improvements on the peephole design that avoid the placement of return SWAPs would involve extensive modifications to Qiskit’s compilation stages or its underlying passes. In Section 4, we will explore such a modification of SABRE that accounts for the only-the-fly changes in qubit layout associated with inter-chiplet routing as an element of our design, in addition to numerous other components to optimize compilation for modular quantum architectures.
4. The Design of SEQC

### v3

@@ Abstract.
As quantum computing technology matures, industry is adopting modular quantum architectures to keep quantum scaling on the projected path and meet performance targets. However, the complexity of chiplet-based quantum devices, coupled with their growing size, presents an imminent scalability challenge for quantum compilation. Contemporary compilation methods are not well-suited to chiplet architectures — in particular, existing qubit allocation methods are often unable to contend with inter-chiplet links, which don’t necessarily support a universal basis gate set. Furthermore, existing methods of logical-to-physical qubit placement, swap insertion (routing), unitary synthesis, and/or optimization, are typically not designed for qubit links of significantly varying latency or fidelity. In this work, we propose SEQC, a hierarchical parallelized compilation pipeline optimized for chiplet-based quantum systems, including several novel methods for qubit placement, qubit routing, and circuit optimization. SEQC attains a $9.3\%$ average increase in circuit fidelity (up to $49.99\%$). Additionally, owing to its ability to parallelize compilation, SEQC achieves 3.27×3.27$\times$3.27 × faster compilation on average (up to 6.74×6.74$\times$6.74 ×) over a chiplet-unaware Qiskit baseline.
@@ 1. Introduction
We are already observing signs of this shift. Recent developments in quantum chip linking, including flip-chip architectures (gold2021entanglement) and low-loss coaxial cables (niu2023low), suggest that modular designs are the most viable for scaling (smith_2022_chiplets). Many contemporary or upcoming leading quantum systems adopt chiplet-based modular quantum processors, for example using high-bandwidth quantum links between nearest-neighbor quantum chiplets (e.g., carrier-chip couplers in Rigetti Aspen-M (Rigetti_AspenM; gold2021entanglement), or m-couplers in IBM Crossbill (gambetta_2024_IBMcouplers; Mandelbaum_2024_IBMCouplers)), or lower-fidelity, lower-bandwidth, but longer-distance flexible coupling of discrete chips (e.g., l-couplers in IBM Flamingo (gambetta_2024_IBMcouplers; Mandelbaum_2024_IBMCouplers)), or a combination of the above (e.g., c-, m-, and l-couplers in IBM Starling (Mandelbaum_2024_IBMCouplers)), or multi-chip connectivity through tunable couplers and routing chips (Field_2024_RigettiModular). Most major companies have set their sights at modular designs to meet quantum scaling targets in practical ways, and modular quantum processors dominate their latest roadmaps (IBM_2023_Roadmap; IonQ_2024_Roadmap; Rigetti_2024_Roadmap).
@@
Unfortunately, current compilation infrastructure does not support quantum interconnect heterogeneity. The mapping of a quantum program’s logical quantum gates into the native gates supported by the underlying quantum processor, and the mapping of logical qubits to the physical qubits of the quantum hardware, largely determine the number and type of native quantum gates required for the computation, the circuit depth, and its execution time—long programs may not complete successfully as qubits decohere. Different mappings can result in vastly different compiled quantum circuits, with diverse characteristics along all these axes. These details of efficiency can make or break an algorithm in today’s noisy intermediate-scale quantum systems (NISQ) era, so quantum compilers aggressively optimize for all of them. To make matters more challenging, each qubit, coupler, and gate have diverse error profiles that are highly variable both spatially and temporally (Tannu_2019_ErrorVariability; Murali_2019_NoiseMapping; dasgupta2021stability). Unlike classical compilation that is done only once for a given architecture, quantum programs must be recompiled every time before execution [^1] The term “execution” here includes all shots required to produce a final result for the user. Different executions of the same circuit can span qubit calibration windows., as the ideal physical qubits to target change between runs. So, quantum compilers today are faced with the daunting task of optimizing quantum programs across multiple dimensions with often conflicting demands, in a continuously changing environment. Coupled with the hardness of synthesizing unitaries (which is exponential in the number of qubits) and the need for quick and frequent compilation, modern quantum compilers have no option other than to rely on heuristics to perform the task. These heuristics result in a compilation complexity that is quadratic to the number of qubits in the quantum processor (experimentally observed in Figure 12).
@@
Compilation in classical systems is typically performed independently and in parallel for each source file, producing one object file per source. The individual object files are then linked together to construct the executable. We propose a compilation framework for modular quantum processors that works in a similar fashion: it stratifies, i.e., splits, the source quantum circuit into subcircuits small enough to fit in each chiplet, and maps subcircuits to chiplets, and then in parallel elaborates each subcircuit and compiles it for its target chiplet. This Stratify-Elaborate Quantum Compiler (SEQC) stratifies a source program only once for a given chiplet architecture, and performs only the elaboration step recurrently before each execution.
As the qubit capacity $n$ of quantum processors grows exponentially, the $O(n^{2})$ compilation latency of contemporary compilers rises even faster. SEQC replaces the recurrent $O(n^{2})$ compilation step with several parallel $O(k^{2})$ elaboration steps for $k$-qubit chiplets. The stratification step is $O(n^{2})$, so the end-to-end complexity remains the same, but stratification is performed only once. The recurrent compilation in SEQC is only $O(k^{2})$, and as the number of qubits $k$ per chiplet is expected to remain relatively stable, or grow at a much slower pace in the foreseeable future (IBM_2023_Roadmap), SEQC’s recurrent compilation latency will barely grow as processors scale.
Additionally, as SEQC is cognizant of hardware modularity, it can stratify the source quantum circuit into subcircuits and map them to chiplets to minimize inter-chiplet communication. As we show in this paper, SEQC produces circuits with shorter execution times and significantly fewer inter-chiplet gates compared to today’s stock compilers, leading to higher execution fidelity. More importantly, as the number of qubits in a processor grows, SEQC achieves even higher performance in these figures of merit.
@@
  • We design and implement in SEQC several novel methods for qubit placement, qubit routing, and circuit optimization. SEQC compiles circuits with $9.3\%$ higher circuit fidelity on average (up to $49.99\%$), while consistently achieving 3.27×3.27$\times$3.27 × faster compilation time on average (up to 6.74×6.74$\times$6.74 ×) compared to a chiplet-unaware Qiskit baseline.
@@ 2. Background and Motivation
(a) Example 2-chiplet architecture
(b) Input circuit from qubit allocation
@@
Qubit allocation addresses the topology constraints of quantum devices, where physical qubits on a device have limited connectivity to other qubits on the device. It is critical that after this stage, which often requires the modification or addition of circuit gates, the resulting optimized, technology-dependent circuit is functionally equivalent to the original, technology-independent algorithm (smith2019quantum). The qubit allocation problem has been previously shown to be NP-complete (qubit_allocation). Thus, algorithms often adopt various strategies to make the problem more tractable. One common method is to further divide qubit allocation into two substages: layout and routing. Logical-to-physical qubit layout, also known as qubit placement, involves deriving an injective map from the virtual qubits of the quantum circuit to the physical qubits of the quantum device (qubit_routing; sabre_swap). In other words, the qubit layout stage decides the initial position of logical qubits on the device. In contrast, qubit routing operates from an initial qubit layout and inserts SWAP operations to align a quantum circuit’s gates to a given qubit topology. Consider, for example, the state-of-the-art in qubit allocation: the SABRE heuristic search algorithm (sabre_swap). In SABRE, the distinction of the layout and routing stages motivates an iterative method for solving qubit allocation. Namely, an initial layout is used to perform qubit routing, which is used to generate a new initial layout.
@@
Because qubit routing remains NP-hard (qubit_routing), another method of simplifying the problem is to assume qubit links are roughly equivalent, particularly with respect to average fidelity and two-qubit gate duration. This assumption can manifest implicitly when a given device topology is represented as an unweighted, undirected graph. For monolithic architectures, one could debate the marginal tradeoff between performance and solution quality. However, for modular architectures, the cross-chiplet connections are far worse compared to intra-chiplet connections (smith_2022_chiplets). Thus, we could expect the solutions generated from the simplified model to be worse than a more physically accurate model, for example, in circumstances where the shortest path in terms of the quantity or depth of inserted SWAPs diverges from the shortest path in terms of fidelity, gate duration, or some combination thereof.
Basis Translation
@@
Basis translation converts the gates of an input quantum circuit into gates that are physically supported on hardware. Monolithic quantum architectures typically support a so-called universal basis gate set, where any unitary quantum operation can be constructed using solely the gates in the set (mike_and_ike). Moreover, it has been shown that basis translation can be performed efficiently given the target gate set is universal, and the quantum circuit contains a fixed, finite number of qubits (Solovay–Kitaev). Unfortunately, modular quantum architectures depend upon specialized links to connect otherwise-independent chiplets. In the general case, the cross-chiplet links can only be used for data movement (such as SWAPs) and not computation, meaning these links may not support universal operations (niu2023low). Thus, compilers that depend on the expectation of universality among qubit links (qiskit_basis_translation_errors), and do not expect these heterogenous, non-universal basis gate sets, can generate incompatible outputs when targeting modular architectures.
Optimization
@@
Quantum circuit optimization seeks to prune extraneous operations by finding opportunities to combine, eliminate, and/or parallelize quantum gates (optimization1). Through commutation and cancellation identities, it is possible to search for optimization opportunities by iteratively transforming the quantum circuit without affecting the underlying operation (optimization1). However, in general, optimization is highly time and resource intensive, with the problem scaling exponentially with respect to number of qubits and/or circuit depth (qfactor). As a result, partitioning, the division of a quantum circuit into blocks of quantum gates, is a common tactic among optimization techniques (qfactor; qgo; bqskit). One major consideration in this process is demarcation, i.e., where and when to draw the line between blocks. Namely, one forfeits all optimization opportunities along boundaries of demarcation, so they must be carefully chosen to minimize the negative impact on the output solution. From this perspective, modular architectures are naturally synergistic with partitioning, since cross-chiplet connections present a physically motivated boundary between operations.
3. Compilation for Modular Quantum Systems
@@
Modular quantum architectures introduce new challenges to quantum compilation that impact the correctness of solutions and the speed of compilation. As discussed previously in §2, our model of modular quantum architectures considers the possibility that inter-chiplet links can be non-universal. This is representative of hardware manufacturing constraints (niu2023low), where only a restricted set of operations, e.g., state transfers or SWAPs, may be supported over long-distance couplers (inter-chiplet links). However, many compiler designs (including designs for modular architectures, see §7) experience errors or crashes when presented with non-universal, heterogenous, basis gate sets. For example, the Qiskit compiler (qiskit2024) is explicit in its compatibility for only universal basis gate sets (qiskit_basis_translation_errors). While it is possible to create a Qiskit Backend that supports heterogeneous basis gate sets, the compiler itself does not take them into account. Specifically, the Qiskit routing stage (zou2024lightsabrelightweightenhancedsabre) (an implementation of SABRE) will freely place gates on incompatible physical links, and there are no methods in the compiler to prevent or fix consequent errors. Moreover, compilers are called to target increasingly larger quantum devices, enabled by modularity. To address the mounting pressure of ever-growing compilation workloads, these compilers must be cognizant of classical hardware limitations and tradeoffs between solution quality and compilation latency.
To ameliorate these issues, we present a family of compilers optimized for modular quantum architectures that support chiplet links with non-universal basis gate sets, and leverage modular compilation strategies. In Section 3.1, we propose a peephole compilation pass that can be inserted into any compiler to add support for non-universal basis gate sets. This would allow any compiler to be compatible with chiplet-based devices without significant changes to their underlying implementations. Next, in Section 3.2, we introduce the stratify-elaborate quantum compiler (SEQC), which advocates for a hierarchical two-stage compilation that produces both better circuits for execution and faster compilation times.
3.1. Chiplet-awareness via Peephole Correction
@@
The design of the peephole correction pass was inspired by the “peephole” optimizations commonly used in classical compilation (dragon_book). Peephole passes operate within a nominal peephole window and substitute its contents with a logically equivalent replacement. Notable examples in the context of quantum compilation include basis translation and elimination of self-adjoint gates. Peephole correction (Figure 1) operates on a circuit generated by qubit allocation, where it detects when non-SWAP gates are placed on non-universal inter-chiplet links (Figure 1(b)), and employs a cyclic qubit rotation, i.e., a perfect shuffle permutation, to move the aforementioned gates to nearby intra-chiplet links (Figure 1(c)). Additional “return” SWAPs are then inserted to return the qubits to their initial locations (Figure 1(d)). Overall, for every inter-chiplet gate in the original circuit generated by SABRE, the peephole pass inserts 4 additional SWAP gates — 2 inter-chiplet and 2 intra-chiplet. Ultimately, this degree of overhead is inevitable when the stock, unmodified compiler is unaware of chiplet-specific device constraints.
@@
For example, skipping the return SWAP operations would initially appear to save $2$ SWAPs, thus halving the cost per inter-chiplet gate. However, this hypothetical design would also no longer be a true peephole, since it introduces a net qubit permutation that propagates to downstream gates in the circuit. From the perspective of modular compilation, this results in serial dependencies that prevent parallel execution of the pass. From a routing perspective, the resultant qubit permutations disrupt the post-routing layout of qubits, leading to destructive routing errors that can only be resolved by an entirely new routing pass. Overall, such a hypothetical pass results in substantially worse output circuits compared to the peephole method.
3.2. The Stratify-Elaborate Quantum Compiler